# Time-dependent Regularized 13-Moment Equations with Onsager Boundary Conditions in the Linear Regime

### v1

@@ Appendix A Tables of coefficients in the inverse-power-law model
$\begin{array}[]{| c | c | c | c | c | c |}\hline\cr\hline\cr\eta&m_{51}&m_{52}&m_{53}&m_{54}&m_{55}\\ 5&-&-&-&-&-\\ 7&3.4282\times 10^{-2}&8.5377\times 10^{-2}&5.6547\times 10^{-9}&7.6406\times 10^{-2}&7.4466\times 10^{-7}\\ 10&5.2206\times 10^{-2}&1.4439\times 10^{-1}&4.0044\times 10^{-8}&1.2791\times 10^{-1}&4.6270\times 10^{-6}\\ 17&6.4303\times 10^{-2}&1.9945\times 10^{-1}&1.3228\times 10^{-7}&1.7479\times 10^{-1}&1.3966\times 10^{-5}\\ \infty&7.3794\times 10^{-2}&2.7955\times 10^{-1}&4.0806\times 10^{-7}&2.3986\times 10^{-1}&4.0020\times 10^{-5}\\ \hline\cr\eta&m_{56}&m_{57}&m_{58}\\ 5&-&-&-\\ 7&1.3161\times 10^{-1}&5.5632\times 10^{-2}&4.7964\times 10^{-1}\\ 10&2.8872\times 10^{-1}&1.2259\times 10^{-1}&8.1033\times 10^{-1}\\ 17&4.7958\times 10^{-1}&2.0482\times 10^{-1}&1.1154\\ \infty&8.1526\times 10^{-1}&3.5174\times 10^{-1}&1.5445\\ \hline\cr\eta&m_{61}&m_{62}&m_{63}&m_{64}&m_{65}\\ 5&\frac{2}{5\sqrt{2\pi}}&\frac{7}{5\sqrt{2\pi}}&\frac{8}{25\sqrt{2\pi}}&\frac{48}{125\sqrt{2\pi}}&0\\ 7&1.5598\times 10^{-1}&6.0144\times 10^{-1}&1.1141\times 10^{-1}&1.4441\times 10^{-1}&7.5718\times 10^{-3}\\ 10&1.5101\times 10^{-1}&6.2395\times 10^{-1}&9.9703\times 10^{-2}&1.3592\times 10^{-1}&1.2112\times 10^{-2}\\ 17&1.4747\times 10^{-1}&6.5107\times 10^{-1}&9.0636\times 10^{-2}&1.2909\times 10^{-1}&1.5815\times 10^{-2}\\ \infty&1.4418\times 10^{-1}&7.0335\times 10^{-1}&8.0389\times 10^{-2}&1.2112\times 10^{-1}&2.0298\times 10^{-2}\\ \hline\cr\eta&m_{66}&m_{67}&m_{68}&m_{69}\\ 5&0&2&0&0\\ 7&2.1284\times 10^{-7}&2.0460&9.0196\times 10^{-8}&3.5825\times 10^{-7}\\ 10&1.6606\times 10^{-6}&2.0707&3.0801\times 10^{-7}&2.8135\times 10^{-6}\\ 17&5.1333\times 10^{-6}&2.1281&6.3694\times 10^{-7}&8.7714\times 10^{-6}\\ \infty&1.5172\times 10^{-5}&2.2756&1.2781\times 10^{-6}&2.6311\times 10^{-5}\\ \hline\cr\eta&m_{71}&m_{81}\\ 5&\frac{1}{2\sqrt{2\pi}}&\frac{1}{2\sqrt{2\pi}}\\ 7&2.0678\times 10^{-1}&2.0678\times 10^{-1}\\ 10&2.0996\times 10^{-1}&2.0996\times 10^{-1}\\ 17&2.1152\times 10^{-1}&2.1152\times 10^{-1}\\ \infty&2.1169\times 10^{-1}&2.1169\times 10^{-1}\\ \hline\cr\hline\cr\end{array}$
@@
Appendix B Computation of the coefficients $\boldsymbol{c_{k}^{\ell,n}}$

### v2

@@ Appendix B Tables of coefficients in the inverse-power-law model
$\begin{array}[]{| c | c | c | c | c | c |}\hline\cr\hline\cr\eta&m_{51}&m_{52}&m_{53}&m_{54}&m_{55}\\ 5&-&-&-&-&-\\ 7&3.4282\times 10^{-2}&8.5377\times 10^{-2}&5.6547\times 10^{-9}&7.6406\times 10^{-2}&7.4466\times 10^{-7}\\ 10&5.2206\times 10^{-2}&1.4439\times 10^{-1}&4.0044\times 10^{-8}&1.2791\times 10^{-1}&4.6270\times 10^{-6}\\ 17&6.4303\times 10^{-2}&1.9945\times 10^{-1}&1.3228\times 10^{-7}&1.7479\times 10^{-1}&1.3966\times 10^{-5}\\ \infty&7.3794\times 10^{-2}&2.7955\times 10^{-1}&4.0806\times 10^{-7}&2.3986\times 10^{-1}&4.0020\times 10^{-5}\\ \hline\cr\eta&m_{56}&m_{57}&m_{58}\\ 5&-&-&-\\ 7&2.7746\times 10^{-2}&5.5632\times 10^{-2}&4.7964\times 10^{-1}\\ 10&6.0867\times 10^{-2}&1.2259\times 10^{-1}&8.1033\times 10^{-1}\\ 17&1.0110\times 10^{-1}&2.0482\times 10^{-1}&1.1154\\ \infty&1.7187\times 10^{-1}&3.5174\times 10^{-1}&1.5445\\ \hline\cr\eta&m_{61}&m_{62}&m_{63}&m_{64}&m_{65}\\ 5&\frac{2}{5\sqrt{2\pi}}&\frac{7}{5\sqrt{2\pi}}&\frac{8}{25\sqrt{2\pi}}&\frac{48}{125\sqrt{2\pi}}&0\\ 7&1.5598\times 10^{-1}&6.0144\times 10^{-1}&1.1141\times 10^{-1}&1.4441\times 10^{-1}&7.5718\times 10^{-3}\\ 10&1.5101\times 10^{-1}&6.2395\times 10^{-1}&9.9703\times 10^{-2}&1.3592\times 10^{-1}&1.2112\times 10^{-2}\\ 17&1.4747\times 10^{-1}&6.5107\times 10^{-1}&9.0636\times 10^{-2}&1.2909\times 10^{-1}&1.5815\times 10^{-2}\\ \infty&1.4418\times 10^{-1}&7.0335\times 10^{-1}&8.0389\times 10^{-2}&1.2112\times 10^{-1}&2.0298\times 10^{-2}\\ \hline\cr\eta&m_{66}&m_{67}&m_{68}&m_{69}\\ 5&0&2&0&0\\ 7&2.1284\times 10^{-7}&2.0460&9.0196\times 10^{-8}&3.5825\times 10^{-7}\\ 10&1.6606\times 10^{-6}&2.0707&3.0801\times 10^{-7}&2.8135\times 10^{-6}\\ 17&5.1333\times 10^{-6}&2.1281&6.3694\times 10^{-7}&8.7714\times 10^{-6}\\ \infty&1.5172\times 10^{-5}&2.2756&1.2781\times 10^{-6}&2.6311\times 10^{-5}\\ \hline\cr\eta&m_{71}&m_{81}\\ 5&\frac{1}{2\sqrt{2\pi}}&\frac{1}{2\sqrt{2\pi}}\\ 7&2.0678\times 10^{-1}&2.0678\times 10^{-1}\\ 10&2.0996\times 10^{-1}&2.0996\times 10^{-1}\\ 17&2.1152\times 10^{-1}&2.1152\times 10^{-1}\\ \infty&2.1169\times 10^{-1}&2.1169\times 10^{-1}\\ \hline\cr\hline\cr\end{array}$
@@
Appendix C Computation of the coefficients $\boldsymbol{\beta}$’s and $\boldsymbol{\gamma}$’s in the asymptotic analysis
This section establishes the relationship between the coefficients $a_{lnn^{\prime}}$ and $\beta$’s and $\gamma$’s in (30)–(34). For any fixed $l$, we consider $a_{lnn^{\prime}}$ as an infinite matrix. We can then define $b_{lnn^{\prime}}^{(n_{0})}$ as the inverse of a submatrix of $a_{lnn^{\prime}}$:
$\sum_{n^{\prime}=n_{0}}^{+\infty}a_{lnn^{\prime}}b_{ln^{\prime}n^{\prime\prime}}^{(n_{0})}=\delta_{nn^{\prime\prime}},\qquad\forall n,n^{\prime\prime}=n_{0},n_{0}+1,\ldots.$
Due to the conservation of mass, momentum and energy, the quantities $b_{0nn^{\prime}}^{(0)}$, $b_{0nn^{\prime}}^{(1)}$ and $b_{1nn^{\prime}}^{(0)}$ do not exist (see (28)). Since the collision operator $\mathcal{L}$ is negative semidefinite and has a kernel of five dimensions, all other coefficients $b_{lnn^{\prime}}^{(n_{0})}$ with nonnegative $l$, $n$, $n^{\prime}$ and $n_{0}$ are well defined. These coefficients have been defined in [10] where R13 equations for steady-state flows are derived.
In practice, the infinite matrices $a_{lnn^{\prime}}$ are truncated and $b_{lnn^{\prime}}^{(n_{0})}$ are obtained by taking the inverse directly. Then the coefficients $\beta$’s and $\gamma$’s in (30)–(34) are given by
$\beta_{1}^{(1),n}=b_{11n}^{(1)},\qquad\beta_{2}^{(1),n}=b_{20n}^{(0)},$
$\gamma^{(2),n}_{0}=\sum_{n^{\prime}=2}^{+\infty}\frac{b_{0nn^{\prime}}^{(2)}(\sqrt{2n^{\prime}+3}b_{11n^{\prime}}^{(1)}-\sqrt{2n^{\prime}}b_{11,n^{\prime}-1}^{(1)})}{b_{111}^{(1)}},$
$\gamma_{t1}^{(1),n}=\sum_{n^{\prime}=2}^{\infty}\frac{b^{(2)}_{1nn^{\prime}}b^{(1)}_{11n^{\prime}}}{b^{(1)}_{111}},\qquad\gamma^{(1),n}_{s1}=\sum_{n^{\prime}=2}^{\infty}\frac{b_{1nn^{\prime}}^{(2)}(\sqrt{2n^{\prime}+5}b_{20n^{\prime}}^{(0)}-\sqrt{2n^{\prime}}b_{20,n^{\prime}-1}^{(0)})}{b_{200}^{(0)}},$
$\gamma_{t2}^{(1),n}=\sum_{n^{\prime}=2}^{\infty}\frac{b^{(1)}_{2nn^{\prime}}b^{(0)}_{20n^{\prime}}}{b^{(0)}_{200}},\qquad\gamma^{(1),n}_{s2}=\frac{2}{5}\sum_{n^{\prime}=1}^{\infty}\frac{b_{2nn^{\prime}}^{(1)}(\sqrt{2n^{\prime}+5}b_{11n^{\prime}}^{(1)}-\sqrt{2(n^{\prime}+1)}b_{11,n^{\prime}+1}^{(1)})}{b_{111}^{(1)}},$
$\gamma^{(2),n}_{3}=\frac{3}{7}\sum_{n^{\prime}=0}^{+\infty}\frac{b_{3nn^{\prime}}^{(0)}}{b_{200}^{(0)}}\left(\sqrt{2n^{\prime}+7}b_{20n^{\prime}}^{(0)}-\sqrt{2(n^{\prime}+1)}b_{20,n^{\prime}+1}^{(0)}\right).$
These results are obtained by asymptotic analysis. The derivation is similar to the steady-state case. Here we omit the details and refer interested readers to [10].
Appendix D Computation of the coefficients $\boldsymbol{c_{k}^{\ell,n}}$